Two charges lie in a line along the x axis. Charge 1 is q1 = 0.65 C and charge 2 is q2 = 2.45 C. They are each a distance of d = 0.021 m from the origin.



What is the distance on the x-axis from the origin at which the electric field will be zero. Give your answer in meters.


1
Expert's answer
2022-01-31T10:35:00-0500

Electric field due to charge "q_1"

"E_2=\frac{kq_1}{r_1^2}"

Electric Field due to charge "q_2"

"E_2=\frac{kq_2}{r^2}"

Net electric field sum of both charge

"E_{net}=E_1-E_2"

"E=\frac{kq_1}{r_1^2}-\frac{kq_2}{r_2^2}=0" P

"r_1=r-x,r_2=r+x"

"\frac{q_1}{(r-x)^2}=\frac{q_2}{(r+x)^2}"


"q_2(r^2-2rx+x^2)=q_1(r^2+2rx+x^2)"

"x^2(q_2-q_1)-2rx(q_1+q_2)-r^2(q_2-q_1)"

"1.8x^2-6.2rx-1.8r^2"


"1.8x^2-6.2\times0.021x-1.8\times0.021"

"1.8x^2-0.1302x-0.0378"

"x=\frac{-b+\sqrt{b^2-4ac}}{2 a}"

"x=\frac{0.1302+\sqrt{0.1302^2+4\times1.8\times0.0378}}{2 \times1.8}" x=0.185m